A lump-sum loan of $15,000 is needed by
Chandra to pay for college expenses. She has obtained
small consumer loans with 10% interest per year in
the past to help pay for college. But her father has
advised her to apply for a PLUS student loan charging
only 7% interest per year. If the loan will be repaid in
full in 5 years, what is the difference in total interest
accumulated by these two types of student loans?
1
Expert's answer
2020-03-31T09:04:54-0400

Total interest = "FV = PV*(1+i)^n"

1) i=10%;

PV = 15,000*(1.1)^5= $24,157.65

2)i=7%;

PV = 15,000*(1.07)^5= $21,038.27.

So, the difference in total interest accumulated is $24,157.65-$21,038.27 = $3,119.38